A random sample of 400 tins of vegetable oil, labelled ""15 kgs net weight"' gave a mean weight of 14.25 kgs with a standard deviation of 1.50 kg. Do we reject the hypothesis of net weight of 15 kgs per tin on the basis of this sample, at 5% level of significance?


1
Expert's answer
2021-02-19T13:52:22-0500

We have that

"\mu = 15"

"n = 400"

"\bar x=14.25"

"s=1.50"

"\alpha=0.05"


"H_0:\mu = 15"

"H_a:\mu \ne15"

The hypothesis test is two-tailed.

Since the population standard deviation is unknown we use the t-test.

The critical value for 5% significance level and 399 df is "\pm1.9659"

(degrees of freedom df = n – 1 = 400 – 1 = 399)

The critical region contains all values smaller than –1.9659 and larger than 1.9659

Test statistic:


"t=\frac{\bar x - \mu}{\frac{s}{\sqrt n}}=\frac{14.25- 15}{\frac{1.5}{\sqrt {400}}}=-10"


Since –10 < –1.9659 thus t falls in the rejection region we reject the null hypothesis.

At the 5% significance level the data do provide sufficient evidence to reject the hypothesis. We are 95% confident to conclude that the mean weight is not 15kgs per tin on the basis of this sample.